# Two-Stage Learning of Stabilizing Neural Controllers via Zubov Sampling and Iterative Domain Expansion

### v1

@@ 4.3 Ablation Studies
Table 3: Example verification time for each solver across our benchmark systems with formal verification. For 2D quadrotor, it is marked with a star due to only forward invariance of the set $V^{\leq 0.2}$ is verified by involving theorem 3.1 with $c_{1}=0.2,c_{2}=0.2+10^{-5}$.
@@
$\alpha,\!\beta$-CROWN 3.94 s 3.00 s 3.64 s 3.94 s 3.77 s 3.67 s 76443 s 209504 s
@@ 5 Conclusion
In this work we propose a two stage training pipeline with novel training domain and training data selection for stabilizing controller with large region of attraction. Through numerical experiments, we demonstrate that our training pipeline can achieve ROA volumes that are $5-1.5\cdot 10^{5}$ bigger than the baselines. We also enable the state-of-the-art verifier $\alpha,\!\beta$-CROWN with the capability of doing bound propagation through many Jacobian operators and introduce novel verification scheme to avoid expensive bisection. We demonstrated that our verification can yield a $40-10000$ times speedup compared to the previously most commonly used verifier dReal.
Although our method achieves promising empirical performance regarding the ROA volume and dimensionality compared to existing baselines, formal verification is still challenging for higher dimensional systems. Currently, we are only able to formally verify 6D systems for forward invariance, but we are interested in exploring the improvement on the verifier side to better aid the formal verification in our context.
@@ A.1 Domain Update Algorithm
1:Controller $u_{\theta}$, Lyapunov Function $V_{\theta}$, Domain $\Omega$, Integration Time $T$ and Discretizations Timesteps $dt$, Number of Trajectories $N_{T}$, Lyapunov Upper Threshold $c$, Domain Expansion Factor $\beta$
2:$x_{\text{init}}\leftarrow$ $N_{T}$ samples from the sublevel set $0<V_{\theta}<c$triangleright Data from (11)
3:$\mathcal{C}\leftarrow$ Set of indices $j$ where trajectory $\phi(t,x_{\text{init}}^{j})$ converges
@@
7:     $\Omega_{d}$leftarrow $[d_{\min},d_{\max}]$
8:$\Omega\leftarrow$ $\beta\times\Omega$triangleright To handle monotonic trajectories.
@@ A.2 CEGIS Dataset
The CEGIS dataset from line $12$ of algorithm 1 will not grow without limitations. In practice, we fix a maximal size $k$ allowed for this dataset. After each iteration, we sort the dataset in descending order according to the violation of Lyapunov condition, and only keep the top $k$ samples.
A.3 Full Loss Function
As we have discussed in the main text, the training loss is simply an adaptation of the Zubov theorem. The goal is to guide $V_{\theta}$ to behave more like the Zubov function through training. We adopt a similar loss function as in [33] and modify the data loss using the novel derivation (13). Moreover, to ensure that the second CEGIS stage is numerically stable, we minimize $V_{\theta}(0)$ only with a hinge-type loss. Let $c$ be a small number, the overall loss function is then a weighted combination of the following parts:
$L_{\text{zero}}(\theta)$ $=\text{ReLU}(V_{\theta}^{2}(0)-c)$ (19)
$L_{\text{pde}}(\theta)$ $=\frac{1}{k_{1}}\sum_{i=1}^{k_{1}}(\nabla V_{\theta}(x_{i})^{\top}f(x_{i},\text{detach}(u(x_{i})))+a(1+V_{\theta}(x_{i}))(1-V_{\theta}(x_{i}))\|x_{i}\|^{p})^{2}$ (20)
[…]
$L_{\text{controller}}(\theta)$ $=\frac{1}{k_{1}}\sum_{i=1}^{k_{1}}\text{detach}(\nabla V_{\theta}(x_{i}))^{\top}f(x_{i},u(x_{i}))$ (22)
$L_{\text{boundary}}(\theta)$ $=\frac{1}{k_{3}}\sum_{i=1}^{k_{3}}(V_{\theta}(y_{i})-1)^{2}$ (23)
Here $x_{i}$ are training data from our novel data selection (11) and (12). $y_{i}$ are boundary points from $\partial(2\Omega)$ and this boundary loss would help during the initial phase of learning by ensuring that the controller does not drive the trajectory to escape from at least two times the training domain. To ensure a balanced weight of the loss term and ease the hyperparameter tuning, we adopt loss weights that are learnable using the techniques presented in [15].
@@ A.4 Hyperparameters
Compared to previous works, our training framework significantly reduces the difficulty of hyperparameter tuning when adapting to new systems. For example, the training domain now does not require too much tuning, and the training almost always works by starting from a small domain. However, several key hyperparameters still require careful adjustment.
Firstly, the weights for each loss term must be chosen with care. In practice, we have found that the boundary loss and the zero loss benefit from higher weights compared to the other three loss terms. Conversely, assigning a relatively small weight to the PDE residual loss tends to yield better results. We hypothesize that this is because an overly dominant PDE loss introduces a complex optimization landscape, which can impede the learning of other critical components. Secondly, the integration involved in the data loss must be approximated using a numerical scheme. Both the discretization timestep and the number of simulation steps play an important role in ensuring stable and efficient learning. We found that a simple forward Euler scheme, with a small timestep such as $dt=0.001\,\mathrm{s}$ and approximately $50$ simulation steps, is often sufficient. In this context, as long as $dt$ is not excessively small, a smaller $dt$ generally performs better than a larger one, since the dynamics can exhibit stiffness when the controller is only partially learned. Finally, during trajectory simulations used for domain updates, a smaller $dt$ and longer steps are recommended. Using a large timestep in this setting can cause otherwise convergent trajectories to be recognized as non-convergent.
@@ Appendix B Verification Details
B.2 Linear Relaxation of $\frac{\mathrm{d}}{\mathrm{d}x}\tanh(x)$ and $\frac{\mathrm{d}}{\mathrm{d}x}\mathrm{sigmoid}(x)$
@@
where parameters $\alpha^{L}$, $\beta^{L}$, $\alpha^{U}$, $\beta^{U}$ are dependent on $l$ and $u$. In our following discussion, for a "bell-shaped" functions $\sigma(x)$, we denote ±$\pm z_{\sigma}$ as the two inflection points (i.e. the solution to $\sigma^{\prime\prime}(x)=0$), and we assume that $l+u\geq 0$ without loss of generality. In many cases, the linear bound is chosen to be a tangent line of $\sigma$, and thus is solely determined by the tangent point $d^{L}$ or $d^{U}$.
@@
  • For the upper bound, let $d_{u}$ be the tangent point of the tangent line that passes through $(u,\sigma(u))$. If $l\geq d_{u}$, the upper bound is chosen to be the line directly connecting two endpoints. Otherwise, $d^{U}$ can be any point between 0 and $d_{u}$ (e.g. linearly interpolated given $l$ ranging from $-u$ to $d_{u}$).
Figure 6: Illustration of linear relaxations of a "bell-shaped" function (sigmoid). (a) $-z_{\sigma}\leq l\leq u\leq z_{\sigma}$. The segment to be bounded is concave. (b) $z_{\sigma}\leq l\leq u$. The segment to be bounded is convex. (c) $l<z_{\sigma}<u$. The segment is bounded similarly to an "S-shaped" function.
B.3 Verification Algorithm
Although a CEGIS procedure has been applied during training, small counterexamples can still exist. Previous works use bisection to identify the optimal boundaries of the region where the Lyapunov condition can be verified. This requires test running the entire verification procedure every time a new candidate boundary is set, which is very inefficient. In contrast, our work, thanks to the formulation of the region to verify as a levelset defined by $c_{1}$ and $c_{2}$ in 17, and the input-split branch-and-bound pipeline, can adaptively adjust the values of $c_{1}$ and $c_{2}$ in a single verification run, avoiding the tedious and time-consuming bisection process.
@@ B.3 Verification Algorithm
Algorithm 3 Adaptive Branch-and-Bound Verification with Levelset Adjustment
@@
2:$\mathcal{S}\leftarrow\text{stack}([\Omega])$triangleright stack of unverified subdomains
@@
4:     $\mathcal{B}\leftarrow\textsc{PopBatch}(\mathcal{S},B)$triangleright pop up to $B$ subdomains
5:     $\mathcal{U}\leftarrow\emptyset$triangleright unverified subdomains in this batch
@@
8:         if condition 17 is not verified to satisfy then $\triangleright$ $c_{1}$ and $c_{2}$ are used here
@@
11:         $D_{\mathrm{worst}}\leftarrow\textsc{WorstDomain}(\mathcal{U})$triangleright greatest positive upper bound of $\dot{V}(x)$
@@
25:         $c_{1}\leftarrow v+\epsilon$triangleright counterexample near lower bound
@@
27:         $c_{2}\leftarrow v-\epsilon$triangleright counterexample near upper bound
@@ C.1 Continuous Time Baselines
In the main text, we discussed the limitations of the two continuous-time baselines [4, 33]. A critical condition for theoretical guarantees is that the controller satisfies $u(x^{*})=u^{*}$; otherwise, $x^{*}$ is no longer an equilibrium point, rendering the Lyapunov-based guarantee vacuous. In NLC [4], we observed that the provided code mistakenly uses a bias-free controller for both the inverted pendulum and the path-following dynamics. While this is correct for the inverted pendulum, where $u^{*}=0$, it is incorrect for the path-following dynamics, as the chosen parametrization still enforces $u(0)=0$, which does not hold in this case. This exact issue has also been noted in an open issue on NLC’s official GitHub repository: https://github.com/YaChienChang/Neural-Lyapunov-Control/issues/14. For the method proposed by Wang et al. [33], we found no special treatment to enforce this condition. In fact, when testing the publicly released checkpoints from their GitHub repository, we found that none of them satisfied $u(x^{*})=u^{*}$. For example, in the Double Integrator system, $u(0)=0.3733$, and in the Van der Pol system, $u(0)=-0.0633$, despite both systems having $u^{*}=0$ as their correct equilibrium input.
@@ C.2 Dynamical Systems
$\dot{\theta_{1}}=\theta_{2},\quad\dot{\theta_{2}}=-\frac{\beta}{ml^{2}}\theta_{2}+\frac{g}{l}\sin(\theta_{1})+\frac{1}{ml^{2}}u.$
@@
$\dot{d}_{e}=v\sin(\theta_{e}),\quad\dot{\theta}_{e}$ $=\frac{v}{l}u-\frac{\cos(\theta_{e})}{\frac{r}{v}-\sin(\theta_{e})}.$
@@
[…]
@@
[…]
@@ C.3 Examples of Domain Update
We present some examples on the effect of domain updates and demonstrate its effectiveness. The initial domain for 2D Quadrotor is picked by 0.4 multiplied the original domain from [40, 26], and the other ones are designed to be small but overall are picked quite arbitrarily without much tuning. We can notice that our domain update mechanism is able to enlarge the training domain and thus the ROA by a large amount. See table 4 for details.
@@
Double Integrator ±$\pm[1,1]$$\pm[27.6,9.6]$
Van der Pol ±$\pm[1,1]$$\pm[6,8.5]$
Pendulum (small torque) ±$\pm[1,2]$$\pm[27.6,69.6]$
Pendulum (large torque) ±$\pm[1,2]$$\pm[30,120]$
Path Tracking (small torque) ±$\pm[2,2]$$\pm[9.6,7.2]$
Path Tracking (large torque) ±$\pm[2,2]$$\pm[12,12]$
Cartpole ±$\pm[0.4,0.4,0.4,0.4]$$\pm[4.8,3.6,13.2,13.2]$
2D Quadrotor ±$\pm[0.3,0.3,0.2\pi,1.6,1.6,1.2]$$\pm[12,13.2,12,19.2,20.4,88.8]$
PVTOL ±$\pm[0.4,0.4,0.4,0.4,0.4,0.4]$$\pm[3.6,3.6,3.6,6,8.4,33.6]$
Ducted Fan ±$\pm[0.4,0.4,0.4,0.4,0.4,0.4]$$\pm[6,7.2,10.8,4.8,3.6,25.2]$

### v2

@@ 2-D Quadrotor Two-stage with Random sampling 0 0 0
Table 3: Example verification time for each solver across our benchmark systems with formal verification. For 2D quadrotor, it is marked with a star due to only forward invariance of the set $V^{\leq 0.2}$ is verified by involving theorem 3.1 with $c_{1}=0.2,c_{2}=0.21$.
@@
$\alpha,\!\beta$-CROWN 3.94 s 3.00 s 3.64 s 3.94 s 3.77 s 3.67 s 76443 s 104614 s
@@ 5 Conclusion
In this work, we propose a two stage training pipeline with novel training domain and training data selection for obtaining stabilizing controller with large region of attraction. Through extensive numerical experiments across various dimensions, we demonstrate that our training pipeline can achieve ROA volumes that are $5-1.5\cdot 10^{5}$ bigger than the baselines. We also extend the state-of-the-art verifier $\alpha,\!\beta$-CROWN with the capability of performing bound propagation through many Jacobian operators and introduce novel verification schemes to avoid expensive bisection. We demonstrated that our verification can yield a 4010,00040-10{,}000 times speedup compared to the previously most commonly used verifier dReal. Although our method achieves promising empirical performance regarding the ROA volume and dimensionality compared to existing baselines, formal verification is still challenging for higher dimensional systems. As future works, it will be interesting to explore more verifier-friendly training strategies and to further tailor and improve the verifier itself for this specific problem. Moreover, it would be interesting to extend the framework to more general control system settings, such as discrete-time or hybrid systems, slowly time-varying systems, or to synthesize and verify robust controller for systems with perturbations.
@@ A.1 Domain Update Algorithm
1:Controller $u_{\theta}$, Lyapunov Function $V_{\theta}$, Domain $\Omega$, Integration Time $T$ and Discretizations Timesteps dtdt, Number of Trajectories $N_{T}$, Lyapunov Upper Threshold $c$, Domain Expansion Factor $\beta$
[…]
@@
7:  $\Omega_{d}$ $\leftarrow$ $d_{\min},d_{\max}]$
[…]
@@ A.2 CEGIS Dataset
The dataset from line $12$ of algorithm 1 will not grow without limitations. In practice, we fix a maximal size $k$ allowed for this dataset. After each iteration, we sort the dataset in descending order according to the violation of Lyapunov condition, and only keep the top $k$ samples.
A.3 Loss Function
As we have discussed in the main text, the training loss is simply an adaptation of the Zubov theorem. The goal is to guide $V_{\theta}$ to behave more like the Zubov function through training. We adopt a similar loss function as in [42] and modify the data loss using the novel derivation (10). Moreover, to ensure that the second CEGIS stage is numerically stable, we minimize $V_{\theta}(0)$ only with a hinge-type loss. In the loss functions, $x_{i}$ are training data from our novel data selection (8) and (9). $y_{i}$ are boundary points from $\partial(2\Omega)$ and this boundary loss would help during the initial phase of learning by ensuring that the controller does not drive the trajectory to escape from at least two times the training domain. To ensure a balanced weight of the loss term and ease the hyperparameter tuning, we adopt loss weights that are learnable using the techniques presented in [20].
@@ A.4 Hyperparameters
Compared to previous works, our training framework significantly reduces the difficulty of hyperparameter tuning when adapting to new systems. For example, the training domain now does not require too much tuning, and the training almost always works when starting from a small domain. However, several key hyperparameters still require careful adjustment.
Firstly, the weights for each loss term must be chosen with care. In practice, we have found that the boundary loss and the zero loss benefit from higher weights compared to the other three loss terms. Conversely, assigning a relatively small weight to the PDE residual loss tends to yield better results. We hypothesize that this is because an overly dominant PDE loss introduces a complex optimization landscape, which can impede the learning of other critical components. Secondly, the integration involved in the data loss must be approximated using a numerical scheme. Both the discretization timestep and the number of simulation steps play an important role in ensuring stable and efficient learning. We found that a simple forward Euler scheme, with a small timestep such as $dt=0.001\,\mathrm{s}$ and approximately $50$ simulation steps, is often sufficient. In this context, as long as $dt$ is not excessively small, a smaller $dt$ generally performs better than a larger one, since the dynamics can exhibit stiffness when the controller is only partially learned. Moreover, during trajectory simulations used for domain updates, a smaller $dt$ and longer steps are recommended. Using a large timestep in this setting can cause otherwise convergent trajectories to be recognized as non-convergent. Finally, in the training data selection, the choice of $c$ is crucial. We typically fix it to be very close to 1, like $c=0.95$. However, as the dimension of the system goes up, the Zubov equation might not be fitted well, so we reduce $c$ to be around $0.9$ for better training stability.
@@ Appendix B Verification Details
B.2 The $\alpha,\!\beta$-CROWN verifier
$\alpha,\!\beta$-CROWN has been developed to be the state-of-the-art neural network verifier, where formally, the toolbox aims to certify
$F(x)\leq 0\text{ for all }x\in\mathcal{B}$ (22)
with $F$ being a general computation graph and $B$ being a hyper-rectangle.
High Level Intuition: $\alpha,\!\beta$-CROWN couples fast symbolic linear bound propagation with input-space branch-and-bound. For a box $\mathcal{B}$, it constructs an affine upper bound of $F$ and provide certification based on the maximum value of this affine bound. If inconclusive, it splits $\mathcal{B}$ into sub-boxes and repeats. As boxes shrink, local linear relaxations tighten and nonlinearities stabilize, so on small sub-boxes the affine bound is often already exact. The full process is very efficient, as each bound computation essentially costs a backward pass; thus many sub-boxes can be processed in parallel (tens of thousands), enabling rapid pruning of large regions and certification with only coarse bounds on tiny boxes. Moreover, thanks to the tensorized nature of bound propagation and the flexibility of the branch-and-bound procedure, $\alpha,\!\beta$-CROWN can naturally handle general specifications with "or" and "and". In principle, $\alpha,\!\beta$-CROWN can handle all cases a traditional SMT solver, such as dReal, can handle, and in a much more scalable fashion. We refer the readers to the following papers for more details [52, 45, 45, 46, 43, 51].
For our application, since the condition in theorem 3.1 can be effectively rewritten as
$(\nabla V(x)\cdot g(x,u(x))<0)$ $\vee(V(x)>c_{2})\vee(V(x)<c_{2})$ (23)
$(g(x,u(x))\cdot\vec{n}(x)<0)$ $\vee(x\not\in\partial\mathcal{B})\vee(V(x)>c_{2}),$ (24)
and each term is a computation graph, $\alpha,\!\beta$-CROWN can efficiently verify this condition.
B.3 Linear Relaxation of $\frac{\mathrm{d}}{\mathrm{d}x}\tanh(x)$ and $\frac{\mathrm{d}}{\mathrm{d}x}\mathrm{sigmoid}(x)$
@@ B.2 The $\alpha,\!\beta$-CROWN verifier
where parameters $\alpha^{L}$, $\beta^{L}$, $\alpha^{U}$, $\beta^{U}$ are dependent on $l$ and $u$. In our following discussion, for a "bell-shaped" functions $\sigma(x)$, we denote $\pm z_{\sigma}$ as the two inflection points (i.e. the solution to $\sigma^{\prime\prime}(x)=0$), and we assume that $l+u\geq 0$ without loss of generality. In many cases, the linear bound is chosen to be a tangent line of $\sigma$, and thus is solely determined by the tangent point $d^{L}$ or $d^{U}$.
@@
  • For the upper bound, let $d_{u}$ be the tangent point of the tangent line that passes through $(u,\sigma(u))$. If $l\geq d_{u}$, the upper bound is chosen to be the line directly connecting two endpoints. Otherwise, $d^{U}$ can be any point between 0 and $d_{u}$ (e.g. linearly interpolated given $l$ ranging from u-u to $d_{u}$).
Figure 6: Illustration of linear relaxations of a "bell-shaped" function (sigmoid). (a) $z_{\sigma}\leq l\leq u\leq z_{\sigma}$. The segment to be bounded is concave. (b) $z_{\sigma}\leq l\leq u$. The segment to be bounded is convex. (c) $l<z_{\sigma}<u$. The segment is bounded similarly to an "S-shaped" function.
B.4 Verification Algorithm
Although a CEGIS procedure has been applied during training, small counterexamples can still exist. Previous works use bisection to identify the optimal boundaries of the region where the Lyapunov condition can be verified. This requires test running the entire verification procedure every time a new candidate boundary is set, which is very inefficient. In contrast, our work, thanks to the formulation of the region to verify as a levelset defined by $c_{1}$ and $c_{2}$ in 19, and the input-split branch-and-bound pipeline, can adaptively adjust the values of $c_{1}$ and $c_{2}$ in a single verification run, avoiding the tedious and time-consuming bisection process.
@@ B.4 Verification Algorithm
Algorithm 3 Adaptive Branch‑and‑Bound Verification with Levelset Adjustment
@@
[…]
@@
4:  $\mathcal{B}\leftarrow\textsc{PopBatch}(\mathcal{S},B)$ $\triangleright$ pop up to $B$ subdomains
5:  $\mathcal{U}\leftarrow\emptyset$ $\triangleright$ unverified subdomains in this batch
@@
8:   if condition 19 is not verified to satisfy then $\triangleright$ $c_{1}$ and $c_{2}$ are used here
@@
11:   $D_{\mathrm{worst}}\leftarrow\textsc{WorstDomain}(\mathcal{U})$ $\triangleright$ greatest positive upper bound of $\dot{V}(x)$
@@
25:   $c_{1}\leftarrow v+\epsilon$ $\triangleright$ counterexample near lower bound
@@
27:   $c_{2}\leftarrow v-\epsilon$ $\triangleright$ counterexample near upper bound
@@ C.1 Continuous Time Baselines
In the main text, we discussed the limitations of the continuous-time baselines [4, 12, 42]. A critical condition for theoretical guarantees is that the controller satisfies $u(x^{*})=u^{*}$; otherwise, $x^{*}$ is no longer an equilibrium point, rendering the Lyapunov-based guarantee vacuous. In NLC [4], we observed that the provided code mistakenly uses a bias-free controller for both the inverted pendulum and the path-following dynamics. While this is correct for the inverted pendulum, where $u^{*}=0$, it is incorrect for the path-following dynamics, as the chosen parametrization still enforces $u(0)=0$, which does not hold in this case. This exact issue has also been noted in an open issue on NLC’s official GitHub repository: https://github.com/YaChienChang/Neural-Lyapunov-Control/issues/14. For the method proposed by Wang et al. [42], we found no special treatment to enforce this condition. In fact, when testing the publicly released checkpoints from their GitHub repository, we found that none of them satisfied $u(x^{*})=u^{*}$. For example, in the Double Integrator system, $u(0)=0.3733$, and in the Van der Pol system, u(0)=0.0633u(0)=-0.0633, despite both systems having $u^{*}=0$ as their correct equilibrium input. Moreover, we believe there are some subtle issues in the verification formulation of Fossil 2.0. In [12], which underpins Fossil 2.0, we notice that their Certificate 2 misses the boundary condition, which should supposedly ensure either $\{x:V(x)\leq\beta\}\cap\mathcal{X}$ is empty or any trajectory starting from this intersection will remain in $\mathcal{X}$. We have observed that their method can occasionally return false positives, i.e., a controller and Lyapunov function are claimed to be found while the controller cannot stabilize any trajectory. Therefore, on top of their returned results, we ran an additional check with trajectory simulations, and recorded a negative for a run if the controller does not work.
@@ C.2 Dynamical Systems
$\dot{\theta_{1}}$ $=\theta_{2}$
$\dot{\theta_{2}}$ $-\frac{\beta}{ml^{2}}\theta_{2}+\frac{g}{l}\sin(\theta_{1})+\frac{1}{ml^{2}}u.$
@@
$\dot{d}_{e}$ $=v\sin(\theta_{e})$
$\dot{\theta}_{e}$ $=\frac{v}{l}u-\frac{\cos(\theta_{e})}{\frac{r}{v}-\sin(\theta_{e})}.$
@@
$\ddot{\theta}$ $=\frac{1}{l(m_{c}+m_{p}\sin^{2}\theta)}(-u\cos\theta-m_{p}l\dot{\theta}^{2}\cos\theta\sin\theta+(m_{c}+m_{p})g\sin\theta).$
@@
The hyperparameters are specified with $m=0.486,l=0.225,g=9.81,\mathbf{I}=\mathrm{diag}(I_{x},I_{y},I_{z})=\mathrm{diag}(0.0049,\ 0.0049,\ 0.0088),c=\frac{1.1}{29}\approx 0.03793,u^{*}=\frac{mg}{4}=\frac{0.486\times 9.81}{4}=1.190325$. The torque limit of $u\in\mathbb{R}^{4}_{\geq 0}$ is given by $|u|\leq 3.6$.
@@ C.3 Examples of Domain Update
We present some examples on the effect of the dynamic domain expansion mechanism and demonstrate its effectiveness. The initial domain for 2D Quadrotor is picked by 0.4 multiplied the original domain from [48, 35], and the other ones are designed to be small but overall are picked quite arbitrarily without much tuning. A comparison between the initial training domain and the one after the first stage training is presented in Table 4. We can notice that our domain update mechanism is able to enlarge the training domain and thus the ROA by a large amount, resulting in much bigger and much more less conservative ROA estimation compared to the baselines.
@@
Double Integrator $\pm[1,1]$ $\pm[27.6,9.6]$
Van der Pol $\pm[1,1]$ $\pm[6,8.5]$
Pendulum (small torque) $\pm[1,2]$ $\pm[27.6,69.6]$
Pendulum (large torque) $\pm[1,2]$ $\pm[30,120]$
Path Tracking (small torque) $\pm[2,2]$ $\pm[9.6,7.2]$
Path Tracking (large torque) $\pm[2,2]$ $\pm[12,12]$
Cartpole $\pm[0.4,0.4,0.4,0.4]$ $\pm[4.8,3.6,13.2,13.2]$
2D Quadrotor $\pm[0.3,0.3,0.2\pi,1.6,1.6,1.2]$ $\pm[12,13.2,12,19.2,20.4,88.8]$
PVTOL $\pm[0.4,0.4,0.4,0.4,0.4,0.4]$ $\pm[3.6,3.6,3.6,6,8.4,33.6]$
Ducted Fan $\pm[0.4,0.4,0.4,0.4,0.4,0.4]$ $\pm[6,7.2,10.8,4.8,3.6,25.2]$